OpenStudy (wesdg1978):

For dy/dx of cos xy+x^3=y^3 so far I've got: -y sinx+cos x dy/dx+3x^2-3y^2dy/dx Is this right so far or not?

OpenStudy (amistre64):

cos(xy) + x^3 = y^3 -sin(xy) (x'y+xy') + 3x^2 x' = 3y^2 y' -sin(xy) (y+xy') + 3x^2 = 3y^2 y'

OpenStudy (amistre64):

assuming ive read your inital setup correctly of course

OpenStudy (wesdg1978):

Yeah that looks more like the available answers, but I'm not sure why it's that way. Why is it cos (xy) and not cosx y?

OpenStudy (amistre64):

let -sin(xy) = A A(y+xy') + 3x^2 = 3y^2 y' Ay + Axy' + 3x^2 = 3y^2 y' Ay + 3x^2 = 3y^2 y' - Axy' Ay + 3x^2 = (3y^2 - Ax) y' Ay + 3x^2 ---------- = y' 3y^2 - Ax

OpenStudy (amistre64):

cos(xy) is what im assuming you meant with: cos xy

OpenStudy (amistre64):

grouping symbols would help out alot in clearing these things up

OpenStudy (anonymous):

\[\cos xy+x ^{3}=y ^{3}\] diff.w.r.t x, \[-\sin xy \left( x \frac{ dy }{dx }+y*1 \right)=3y ^{2}\frac{ dy }{ dx }\] \[or -x \sin xy \frac{ dy }{ dx }-y \sin xy=3y ^{2}\frac{ dy }{ dx }\] \[\left( 3y ^{2}+x \sin xy \right)\frac{ dy }{dx }=-y \sin xy\] you can find dy/dx

OpenStudy (wesdg1978):

@surjithayer I think you left out the 3x^2 but I got it anyways, and the rest of it matches what I've got. Thanks for helping!

OpenStudy (anonymous):

yw